you are painting a wall with the length 5 X2 feet and a width 12x feet. There is a rectangular door that measures x feet by 2x feet that will not be painted. What is the area of the wall that is to be painted
Answer by Alan3354(69443)

you are painting a wall with the length 5 X2 feet and a width 12x feet. There is a rectangular door that measures x feet by 2x feet that will not be painted. What is the area of the wall that is to be painted
==========
5X2 = 10
Area = 10*12x = 120x
----
Area of door = x*2x = 2x^2
--> 120x - 2x^2